# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ V-A2 Feature-Projection-based 3D object detection
In contrast to Point-Projection-based methods, Feature-Projection-based 3D object detection methods [246, 250, 232, 445, 446, 129, 241, 449, 377] primarily focus on fusing point cloud features with image features during the feature extraction phase of point clouds. During this fusion process, point cloud features are projected onto corresponding image features, and subsequently, these image and point cloud features are integrated together. This process is achieved by applying a calibration matrix to transform the voxel’s three-dimensional coordinate system into the pixel coordinate system of the image, thereby facilitating the effective fusion of point cloud and image modalities. Specifically, the projection of a three-dimensional point cloud onto the image plane can be articulated as follows:
@@
A quintessential example of the Feature-Projection-based method, ContFuse [246], employs continuous convolution to amalgamate multi-scale convolutional feature maps from each sensor. Within this technique, the projection of the point cloud facilitates the correspondence between the image and the Bird’s Eye View (BEV). In essence, Feature-Projection-based 3D object detection method is accomplished during the point cloud feature extraction phase. Compared to Point-Projection-based methods, they do not perform fusion on the original point cloud but achieve a profound depth feature fusion, resulting in more robust performance. space.
@@ V-A3 Auto-Projection-based 3D object detection
As shown in Fig. 13, a partial image from the KITTI [255] dataset exemplifies that projection inaccuracies persist even in this classic clean dataset. Consequently, the issue of projection errors cannot be completely eliminated by manual calibration, but rather, they can only be mitigated. This is a frequent challenge in practical dataset deployments. Many studies like Point & Feature-Projection-based methods have performed fusion through direct projection without addressing the projection error issue. A few works [106, 233, 234, 384, 243, 378, 247, 450] have sought to mitigate these errors through approaches such as projection offsets and neighboring projections. For instance, Deformable Cross Attention[354] has been employed to learn offsets in the context of already projected data. We have systematically reviewed and synthesized methods that tackle projection errors, designating them as Auto-projection-based 3D object detection methods. As representative works addressing feature alignment, HMFI[243], GraphAlign[233], and GraphAlign++[234] utilize the a priori knowledge of projection calibration matrices to project onto corresponding images for local graph modeling. This approach simulates the intermodal relationships, enabling multimodal 3D object detectors to effectively identify more appropriate alignment relationships, thereby achieving faster and more accurate feature alignment between modalities. AutoAlignV2 [384] focuses on sparse learnable sampling points for cross-modal relational modeling, enhancing calibration error tolerance and significantly accelerating feature aggregation across different modalities. In summary, Auto-Projection-based 3D object detection methods mitigate errors arising from feature alignment by leveraging neighbor relationships or neighbor offset, thereby enhancing robustness in multimodal 3D object detection.
@@ V-A4 Decision-Projection-based 3D object detection
Decision-Projection-based 3D object detection methods [451, 452, 453, 380, 381, 196, 454, 382, 107], as early implementations of multimodal 3D object detection schemes, use projection matrices to align features in Regions of Interest (RoI) or specific results. These methods are primarily focused on the alignment of features in localized areas of interest or specific detection outcomes. Graph-RCNN [107] project the graph node to the location in the camera image and collect the feature vector at that pixel in the camera image through bilinear interpolation. F-PointNet [381] performs detection of the 2D image to determine the class and localization of the object, and for each detected object, the corresponding Point clouds in 3D space is obtained through the conversion matrix of calibrated sensor parameters and 3D space. MV3D [453] employs a transformation of the LiDAR point cloud into Bird’s Eye View (BEV) and Front View (FV) projections for generating proposals. During this process, a specialized 3D proposal network is used to create precise 3D candidate boxes. These 3D proposals are then projected onto feature maps from multiple perspectives to facilitate feature alignment between the two modalities. Differing from MV3D [453] , AVOD [452] streamlines this approach by omitting the FV component and introducing a more refined region proposal mechanism. In summary, Decision-Projection-based 3D object detection methods primarily achieve feature fusion at a high-level through projection, with limited interactivity between heterogeneous modalities. This often leads to the alignment and fusion of erroneous features, resulting in issues of reduced accuracy and robustness.
@@ V-B Non-Projection-based 3D object detection
Non-Projection-based 3D object detection methods achieve fusion without relying on feature alignment, thereby yielding robust feature representations. They circumvent the limitations of camera-to-LiDAR projection, which often reduces the semantic density of camera features and impacts the effectiveness of techniques like Focals Conv [232] and PointPainting [222]. Non-Projection-based methods typically employ cross-attention mechanisms or the construction of a unified space to address the inherent misalignment issues in direct feature projection. These methods are primarily divided into two categories: (1) Query-Learning-based [230, 236, 237, 437, 245, 455] and (2) Unified-feature-based [385, 231, 386, 375, 228, 356, 137, 221, 387, 235, 225, 136, 223]. Query-Learning-based methods entirely negate the need for alignment during the fusion process. Conversely, Unified-feature-based methods, though constructing a unified feature space, do not completely avoid projection; it usually occurs within a single modality context. For example, BEVFusion [231] utilizes LSS [316] for camera-to-BEV projection. This process, taking place before fusion, demonstrates considerable robustness in scenarios with feature misalignment.
@@ V-B1 Query-Learning-based 3D object detection
Query-Learning-based 3D object detection methods, as exemplified by works such as [230, 236, 237, 437, 245, 455], eschew the necessity for projection within the feature fusion process. Instead, they attain feature alignment through cross-attention mechanisms before engaging in the fusion of features. Point cloud features are typically employed as queries, while image features serve as keys and values, facilitating a global feature query to acquire highly robust multimodal features. Furthermore, DeepInteraction [237] incorporates multimodality interaction, wherein point cloud and image features are utilized as distinct queries to enable further feature interaction. In comparison to the exclusive use of point cloud features as queries, the comprehensive incorporation of image features leads to the acquisition of more resilient multimodal features. Overall, Query-Learning-based 3D object detection methods employ a transformer-based structure for feature querying to achieve feature alignment. Ultimately, the multimodal features are integrated into LiDAR-based pipelines, such as CenterPoint[125].
@@ V-B2 Unified-Feature-based 3D object detection
Unified-feature-based 3D object detection methods, represented by works such as [385, 231, 386, 375, 228, 356, 137, 221, 387, 235, 225, 136, 223], generally employ projection before feature fusion, achieving the pre-fusion unification of heterogeneous modalities. In the BEV fusion series, which utilizes LSS for depth estimation [385, 231, 386, 223], the front-view features are transformed into BEV features, followed by the fusion of BEV image and BEV point cloud features. Alternatively, CMT [225] and UniTR [356] employ transformers for tokenization of point clouds and images, constructing an implicit unified space through transformer encoding. CMT [225] utilizes projection in the position encoding process, but entirely avoids dependency on projection relations at the feature learning level. FocalFormer3D [375], FUTR3D [228], and UVTR [136] leverage transformers’ queries to implement schemes similar to DETR3D[88], constructing a unified sparse BEV feature space through queries, thus mitigating the instability introduced by direct projection. VirConv[221], MSMDFusion[387], and SFD[235] construct a unified space through pseudo-point clouds, with the projection occurring before feature learning. The issues introduced by direct projection are addressed through subsequent feature learning. In summary, Unified-feature-based 3D object detection methods [385, 231, 386, 375, 228, 356, 137, 221, 387, 235, 225, 136, 223] currently represent high-precision and robust solutions. Although they incorporate projection matrices, such projection does not occur between multimodal fusion, distinguishing them as Non-Projection-based 3D object detection methods. Unlike Auto-Projection-based 3D object detection approaches, they do not directly address projection error issues but instead opt for unified space construction, considering multiple dimensions for multimodal 3D object detection, thereby obtaining highly robust multimodal features.
@@ V-C Analysis: Accuracy, Latency, Robustness
In the preceding Sections III-D, IV-E, we have conducted a comprehensive analysis of ‘Accuracy, Latency, Robustness’ for camera-only and LiDAR-only approaches. Subsequently, we extend our examination to multimodal 3D object detection methods, employing a similar analytical framework.
@@ V-C1 Accuracy
As shown in Fig.14, we conducted comparative evaluations on both the KITTI test and nuScenes test datasets. The majority of Projection-based 3D object detection methods have predominantly undergone experimentation on the KITTI dataset, with only a minority extending their evaluation to nuScenes. As shown in Fig.14 (a), it is evident that Feature-Projection-based and Auto-Projection-based methods exhibit superior overall performance, while Decision-Projection-based methods, primarily dated prior to 2020, tend to manifest relatively lower Average Precision (AP) metrics. A scant few Non-Projection-based 3D object detection methods, such as CAT-Det, have been experimented with on the KITTI dataset. As shown in Fig.14 (b), the latest methods predominantly belong to the Unified-Feature-based methods, underscoring the suitability of the panoramic camera offered by nuScenes for achieving modality-unifying strategies like BEVFusion[231]. Overall, it is discernible that Non-Projection-based methods present more effective solutions in terms of Accuracy metrics (e.g., AP, mAP, NDS, etc.).
@@ V-C2 Latency
As shown in Fig. 6, we conducted a comparative analysis of mono-modal 3D object detection methods (LiDAR-only and Camera-only) and multimodal 3D object detection on the KITTI and nuScenes datasets, presenting scatter plots for Latency (FPS) and Accuracy metrics (AP, mAP, NDS, etc.). It is noteworthy that, in comparison to mono-modal 3D object detection methods (LiDAR-only and Camera-only), multimodal 3D object detection approaches generally exhibit lower FPS.
As shown in Fig. 6 (e), the results on the KITTI dataset indicate that GraphAlign excels in both AP and FPS metrics. Additionally, LoGoNet [247], Focals Conv [232], and EP-Net [226] demonstrate outstanding performance. Yes, the provided text contains a few errors. As shown in Fig. 6 (f), GraphAlign [233] maintains its position as having the highest FPS, but its NDS performance is suboptimal on the nuScenes dataset. In contrast, UniTR performs exceptionally well in both NDS and FPS metrics. Overall, it can be observed that within Projection-based methods, Auto-Projection-based and Feature-Projection-based methods exhibit superior overall performance, while within Unified-Feature-based methods, the overall performance is more outstanding. In the meticulous evaluation of the KITTI and nuScenes datasets, emphasis is placed on the trade-off between FPS and NDS metrics.
@@ V-C3 Robustness
In the previous sections III-D3 and IV-E3, we analyzed the robustness of mono-modal 3D object detection (Camera-only and LiDAR-only). In this section, based on Tables IV and V, we analyze the robustness of multimodal 3D object detection. From KITTI-C [254] and nuScenes-C [254], it can be seen that multimodal 3D object detection is more robust compared to mono-modal 3D object detection (Camera-only and LiDAR-only), with smaller RCE. In KITTI-C, representative articles LoGoNet [247] for Auto-Projection-based and VirConv [221] for Unified-Feature-based exhibit greater robustness, while EPNet [226] for Point-Projection-based and Focals Conv [232] for Feature-Projection-based show slightly weaker performance. Additionally, in nuScenes-C, among Non-Projection-based methods, FUTR3D [228], TransFusion [230], BEVFusion [231], and DeepInteraction [237] all demonstrate strong robustness.
VI Conclusion

### v2

@@ V-A2 Feature-Projection-based 3D object detection
In contrast to Point-Projection-based methods, Feature-Projection-based 3D object detection methods [432, 122, 110, 419, 327, 60, 118, 420, 326], as shown in Fig. 7 (b), primarily focus on fusing point cloud features with image features during the feature extraction phase of the point clouds. During this fusion process, point cloud features are projected onto corresponding image features, and subsequently, these image and point cloud features are integrated together. This process is achieved by applying a calibration matrix to transform the voxel’s three-dimensional coordinate system into the pixel coordinate system of the image, thereby facilitating the effective fusion of point cloud and image modalities. Specifically, the projection of a three-dimensional point cloud onto the image plane can be articulated as follows:
@@
A quintessential example of the Feature-Projection-based method, ContFuse [432], employs continuous convolution to amalgamate multi-scale convolutional feature maps from each sensor. Within this technique, the projection of the point cloud facilitates the correspondence between the image and the Bird’s Eye View (BEV). In essence, Feature-Projection-based 3D object detection method is accomplished during the point cloud feature extraction phase. Compared to Point-Projection-based methods, they do not perform fusion on the original point cloud but achieve a profound depth feature fusion, resulting in more robust performance.
@@ V-A3 Auto-Projection-based 3D object detection
As shown in Fig. 8, a partial image from the KITTI [133] dataset exemplifies that projection inaccuracies persist even in this classic clean dataset. Consequently, the issue of projection errors cannot be completely eliminated through manual calibration; instead, they can only be mitigated. This is a frequent challenge in practical dataset deployments. Many studies, like Point & Feature-Projection-based methods, have performed fusion through direct projection without addressing the projection error issue. A few works [41, 111, 112, 332, 120, 382, 121, 331], have sought to mitigate these errors through approaches such as projection offsets and neighboring projections. For instance, Deformable Cross Attention[373] has been employed to learn offsets in the context of already projected data. We have systematically reviewed and synthesized methods that tackle projection errors, designating them as Auto-projection-based 3D object detection methods, as shown in Fig. 7 (c). As representative works addressing feature alignment, HMFI[120], GraphAlign[111], and GraphAlign++[112] utilize a priori knowledge of projection calibration matrices to project onto corresponding images for local graph modeling. This approach simulates intermodal relationships, enabling Multi-modal 3D object detectors to effectively identify more appropriate alignment relationships, thereby achieving faster and more accurate feature alignment between modalities. AutoAlignV2 [332] focuses on sparse learnable sampling points for cross-modal relational modeling, enhancing calibration error tolerance and significantly accelerating feature aggregation across different modalities. In summary, Auto-Projection-based 3D object detection methods mitigate errors arising from feature alignment by leveraging neighbor relationships or neighbor offsets, thereby enhancing robustness in Multi-modal 3D object detection.
@@ V-A4 Decision-Projection-based 3D object detection
Decision-Projection-based 3D object detection methods [333, 334, 335, 383, 336, 96, 337, 338, 113], as early implementations of Multi-modal 3D object detection schemes, use projection matrices to align features in Regions of Interest (RoI) or specific results, as shown in Fig. 7 (d). These methods are primarily focused on the alignment of features in localized areas of interest or specific detection outcomes.
Graph-RCNN [113] projects the graph node to the location in the camera image and collects the feature vector at that pixel in the camera image through bilinear interpolation. F-PointNet [336] performs detection on the 2D image to determine the class and localization of the object, and for each detected object, the corresponding point clouds in 3D space are obtained through the conversion matrix of calibrated sensor parameters and 3D space. MV3D [335] employs a transformation of the LiDAR point cloud into Bird’s Eye View (BEV) and Front View (FV) projections for generating proposals. During this process, a specialized 3D proposal network is used to create precise 3D candidate boxes. These 3D proposals are then projected onto feature maps from multiple perspectives to facilitate feature alignment between the two modalities. Differing from MV3D [335], AVOD [334] streamlines this approach by omitting the FV component and introducing a more refined region proposal mechanism. In summary, Decision-Projection-based 3D object detection methods primarily achieve feature fusion at a high level through projection, with limited interaction between heterogeneous modalities. This often leads to the alignment and fusion of erroneous features, resulting in issues of reduced accuracy and robustness.
@@ V-B Non-Projection-based 3D object detection
Non-Projection-based 3D object detection methods achieve fusion without relying on feature alignment, thereby yielding robust feature representations. They circumvent the limitations of camera-to-LiDAR projection, which often reduces the semantic density of camera features and impacts the effectiveness of techniques like Focals Conv [110] and PointPainting [100]. Non-Projection-based methods typically employ cross-attention mechanisms or the construction of a unified space to address the inherent misalignment issues in direct feature projection. These methods are primarily divided into two categories: (1) Query-Learning-based [108, 115, 116, 339, 385, 425] and (2) Unified-feature-based [344, 109, 345, 346, 106, 347, 99, 348, 114, 103, 65, 101]. Query-Learning-based methods entirely negate the need for alignment during the fusion process. Conversely, Unified-Feature-based methods, though constructing a unified feature space, do not completely avoid projection; it usually occurs within a single modality context. For example, BEVFusion [109] utilizes LSS [256] for camera-to-BEV projection. This process, taking place before fusion, demonstrates considerable robustness in scenarios with feature misalignment.
@@ V-B1 Query-Learning-based 3D object detection
Query-Learning-based 3D object detection methods, as exemplified by works such as [108, 115, 116, 339, 385, 425, 433], eschew the necessity for projection within the feature fusion process, as shown in Fig. 7 (e). Instead, they attain feature alignment through cross-attention mechanisms before engaging in the fusion of features. Point cloud features are typically employed as queries, while image features serve as keys and values, facilitating a global feature query to acquire highly robust Multi-modal features. Furthermore, DeepInteraction [116] incorporates multimodality interaction, wherein point cloud and image features are utilized as distinct queries to enable further feature interaction. In comparison to the exclusive use of point cloud features as queries, the comprehensive incorporation of image features leads to the acquisition of more resilient Multi-modal features. Overall, Query-Learning-based 3D object detection methods employ a transformer-based structure for feature querying to achieve feature alignment. Ultimately, the Multi-modal features are integrated into LiDAR-only pipelines, such as CenterPoint[57].
@@ V-B2 Unified-Feature-based 3D object detection
Unified-feature-based 3D object detection methods, represented by works such as [344, 109, 345, 346, 106, 347, 99, 348, 114, 103, 65, 101], generally employ projection before feature fusion, achieving the pre-fusion unification of heterogeneous modalities, as shown in Fig. 7 (f). In the BEV fusion series, which utilizes LSS for depth estimation [344, 109, 345, 101], the front-view features are transformed into BEV features, followed by the fusion of BEV image and BEV point cloud features. Alternatively, CMT [103] and UniTR [347] employ transformers for tokenization of point clouds and images, constructing an implicit unified space through transformer encoding. CMT [103] utilizes projection in the position encoding process, but entirely avoids dependency on projection relations at the feature learning level. FocalFormer3D [346], FUTR3D [106], and UVTR [65] leverage transformers’ queries to implement schemes similar to DETR3D[30], constructing a unified sparse BEV feature space through queries, thus mitigating the instability introduced by direct projection. VirConv[99], MSMDFusion[348], and SFD[114] construct a unified space through pseudo-point clouds, with the projection occurring before feature learning. The issues introduced by direct projection are addressed through subsequent feature learning. In summary, Unified-feature-based 3D object detection methods [344, 109, 345, 346, 106, 347, 99, 348, 114, 103, 65, 101] currently represent high-precision and robust solutions. Although they incorporate projection matrices, such projection does not occur between Multi-modal fusion, distinguishing them as Non-Projection-based 3D object detection methods. Unlike Auto-Projection-based 3D object detection approaches, they do not directly address projection error issues but instead opt for unified space construction, considering multiple dimensions for Multi-modal 3D object detection, thereby obtaining highly robust Multi-modal features.
@@ V-C Analysis: Accuracy, Latency, Robustness
In the preceding Sections III-D, IV-E, we have conducted a comprehensive analysis of ‘Accuracy, Latency, Robustness’ for camera-only and LiDAR-only approaches. Subsequently, we extend our examination to Multi-modal 3D object detection methods, employing a similar analytical framework.
@@ V-C1 Accuracy
As shown in Fig.3 (e) and (f), we conducted comparative evaluations on both the KITTI and nuScenes test datasets. The majority of Projection-based 3D object detection methods have predominantly undergone experimentation on the KITTI dataset, with only a minority extending their evaluation to nuScenes. As shown in Fig.3 (e), it is evident that Feature-Projection-based and Auto-Projection-based methods exhibit superior overall performance, while Decision-Projection-based methods, primarily dated prior to 2020, tend to manifest relatively lower Average Precision (AP) metrics. A scant few Non-Projection-based 3D object detection methods, such as CAT-Det [385], have been experimented with on the KITTI dataset. As shown in Fig.3 (f), the latest methods predominantly belong to the Unified-Feature-based methods, underscoring the suitability of the panoramic camera offered by nuScenes for achieving modality-unifying strategies like BEVFusion[109]. Overall, it is discernible that Non-Projection-based methods present more effective solutions in terms of Accuracy metrics (e.g., AP, mAP, NDS, etc.).
@@ V-C2 Latency
As shown in Table V, we conducted a comparative analysis of mono-modal 3D object detection methods (LiDAR-only and Camera-only) and Multi-modal 3D object detection on the KITTI and nuScenes datasets, presenting scatter plots for Latency (FPS) and Accuracy metrics (AP, mAP, NDS, etc.). It is noteworthy that, in comparison to mono-modal 3D object detection methods (LiDAR-only and Camera-only), Multi-modal 3D object detection approaches generally exhibit lower FPS. The results on the KITTI dataset indicate that GraphAlign excels in both AP and FPS metrics. Additionally, LoGoNet [121], Focals Conv [110], and EP-Net [104] demonstrate outstanding performance. GraphAlign [111] maintains its position as having the highest FPS, but its NDS performance is suboptimal on the nuScenes dataset. In contrast, UniTR performs exceptionally well in both NDS and FPS metrics. Overall, it can be observed that within Projection-based methods, Auto-Projection-based and Feature-Projection-based methods exhibit superior overall performance, while within Unified-Feature-based methods, the overall performance is more outstanding. In the meticulous evaluation of the KITTI and nuScenes datasets, emphasis is placed on the trade-off between FPS and NDS metrics.
@@ V-C3 Robustness
In the previous sections III-D3 and IV-E3, we analyzed the robustness of mono-modal 3D object detection (Camera-only and LiDAR-only). In this section, based on Tables VII and VIII, we analyze the robustness of Multi-modal 3D object detection. From KITTI-C [125] and nuScenes-C [125], it can be seen that Multi-modal 3D object detection is more robust compared to mono-modal 3D object detection (Camera-only and LiDAR-only), with smaller RCE. In KITTI-C, representative articles LoGoNet [121] for Auto-Projection-based and VirConv [99] for Unified-Feature-based exhibit greater robustness, while EPNet [104] for Point-Projection-based and Focals Conv [110] for Feature-Projection-based show slightly weaker performance. Additionally, in nuScenes-C, among Non-Projection-based methods, FUTR3D [106], TransFusion [108], BEVFusion [109], and DeepInteraction [116] all demonstrate strong robustness. It is worth noting that MetaBEV [430] explores the problem of modal loss caused by feature misalignments and sensor failures in BEV features of LiDAR and camera through deformable attention based on BEVFusion [109]. ObjectFusion [117] proposes a novel object-centric fusion to align object-centric features of different modalities. GraphBEV [349] mitigates misalignment issues by matching neighbor depth features through graph matching.
VI Future outlooks
Through reviewing all the literature and analyzing the research trends of the past few years, we make some predictions on the future research direction of 3D object detection from the perspective of robustness.
VI-A 3D Object Detection with Large Models
Inspired by the success of large language models (LLMs) such as ChatGPT [434] and vision foundation models (VFMs) like SAM, many researchers have focused on related research of large models. Compared to conventional methods, a large autonomous driving model based on LLM can mainly solve the following two problems. Firstly, it is the endless corner case problem. LLMs have a common sense ability and may become a new paradigm for solving corner cases in autonomous driving problems. Secondly, the current methods lack intuitive reasoning and provide textual explanations, and LLMs happen to be the best in this direction. It is worth researching how to combine large models with 3D object detection to enhance robustness and generalization and improve the ability of corner cases. However, there is currently limited research on combining large models and 3D object detection. For example, RoboFusion [330] has integrated TransFusion [108] and Focals Conv [110] with VFMs like SAM [435] to enhance its ability in harsh weather conditions. SEAL [436] uses VFMs like SAM [435] to segment different car point cloud sequences and can segment any car point cloud by encouraging spatial and temporal consistency during the representation learning stage. CLIP-BEVFormer [281] combines CLIP [437] and BEVFormer [16], leveraging the universal capabilities of CLIP [437] to enhance generalization on corner cases. VisLED [438] is a language-driven active learning framework for open-set 3D object detection, which utilizes active learning techniques to query various information-rich data samples from unlabeled pools. Almost all existing work is proposed and evaluated on close-range datasets. Although these datasets may be large and diverse, they are still insufficient for real-world applications. In the real world, the generalization and robustness of corner cases are of utmost importance, and 3D object detection with large models is a good starting point for solving open-set 3D object detection. In addition, current 3D object detection algorithms lack interpretability, and LLMs can bring hope for more robust 3D object detection and avoid unexpected situations caused by black box detectors.
VI-B 3D Object Detection in End-to-End Autonomous Driving
UniAD[439] undoubtedly brought another hot topic to the field of autonomous driving after winning the CVPR Best Paper Award: end-to-end autonomous driving. End-to-end autonomous driving is a fully differentiable machine learning system that takes raw sensor input data and other metadata as prior information and directly outputs the control signals or trajectory planning for vehicles [440]. Generally, the autonomous drive system integrates multiple tasks, such as detection, tracking, online mapping, motion prediction, and planning. 3D object detection is closely related to other perception tasks and downstream tasks such as prediction and planning. Therefore, pursuing high accuracy in 3D object detection may not be optimal when considering the autonomous drive system. Although impressive progress has been made in end-to-end research, we believe three areas can be further improved in current 3D object detection for end-to-end autonomous driving. First, 3D object detection can guide more effective multi-modal environmental perception, allowing for better data integration from multi-modal sources. Second, the current inference capabilities of end-to-end autonomous driving are concerning. Third, 3D object detection enhanced with large language models (LLMs) provides stronger explanatory power, leading to enhanced explanatory power for subsequent tasks.
VII Conclusion